OpenStudy (snickers):

I'm trying to figure out how to sketch a gragh witha line of y=-2x+5 and label exactly points don't know what to really do?

OpenStudy (anonymous):

First find the intercepts. Plug in y=0. Plug in x=0.

OpenStudy (anonymous):

Y=0 would give us x=5/2 X=0 would give us y=5

OpenStudy (anonymous):

Now if you have drawn a grid. You can easily put atleast these two points in.

OpenStudy (anonymous):

| | <-----would be y=5 when x=0 ------- | ^ | | | would be x=5/2 when y=0

OpenStudy (anonymous):

Then comes our other piece of information: The slope. In our case it is -2. The negative tells us the line is "falling down". The 2 tells us that for every 1 in the x-direction the y changes by 2.

OpenStudy (anonymous):

You can always try plugging in any X of your choice and the equation will spit back at you the Y that goes with it. So if you put in x=1, y= -2(1) + 5 = 3. You can plot the point (1,3) .

OpenStudy (snickers):

cool thank you it's starting to make more sense now.

OpenStudy (anonymous):

Sketching a curve means plotting ALL of these points! But if we know a few we can just connect the dots and it will give us a good idea what it looks like.

OpenStudy (snickers):

wouldn't x=-5/-2?

OpenStudy (anonymous):

The negatives will cancel out. (-1)*(-1) = 1 or -1/-1 = 1